OpenStudy (anonymous):

solve -3(2x -1) = x + 10 help please

OpenStudy (anonymous):

Expand the brackets and then isolate x.....

OpenStudy (solomonzelman):

1) expand the left side 2) add 6x to both sides 3) subtract 10 from both sides 4) divide both sides by 7

OpenStudy (anonymous):

-6x - 3 = x + 10 first step?

OpenStudy (anonymous):

-3 = 6x + 10 for next step ? @SolomonZelman

OpenStudy (solomonzelman):

you get 7x on the right side, not 6x. b/c x+6x=7x

OpenStudy (solomonzelman):

So after the second step, \(\normalsize\color{red}{ \rm -3=7x+10 }\)

OpenStudy (anonymous):

-13 = 7X

OpenStudy (solomonzelman):

\(\normalsize\color{red}{ \rm 3=7x+10 }\) 3 is positive b/c (-1)*(-3)=3

OpenStudy (solomonzelman):

then \(\normalsize\color{red}{ \rm 3=7x+10 }\) \(\normalsize\color{red}{ \rm -7=7x }\) \(\normalsize\color{red}{ \rm -1=x }\)

OpenStudy (anonymous):

ohhhh that makes sense